# Double Successive Over-Relaxation Q-Learning with an Extension to Deep Reinforcement Learning

### v1

@@ IV Theoretical Analysis
Suppose (A1) holds. Given an MDP defined as in Section II, and $w_{n}$ as in (4). Let every state-action pair be sampled indefinitely. Then, for sample $\{i,a,j,r^{j}_{ia}\}$ the update rule of model-free SOR Q-learning algorithm given by:
@@
The correspondence to Lemma 1 follows from associating $X$ with the set of state-action pairs $(i,a)$, $\beta_{n}(x)$ with $\beta_{n}(i,a)$, and $\Phi_{n}(i,a)$ with $Q_{n}(i,a)-Q^{*}_{w^{*}}(i,a)$, where $Q^{*}_{w^{*}}$ is the fixed point of $U_{w^{*}}$. Let the filtration for this process be defined by, $\mathcal{G}_{n}=\{Q_{0},i_{j},a_{j},\beta_{j},w_{j},\forall j\leq n,n\geq 0\}$. Note that $\mathcal{G}_{n}\subseteq\mathcal{G}_{n+1},\forall n\geq 0.$ The iteration of model-free SORQL is rewritten as follows:
$\Phi_{n+1}(i_{n},a_{n})$ $=(1-\beta_{n}(i_{n},a_{n}))\Phi_{n}(i_{n},a_{n})$
@@
$\quad\left.+(1-w_{n})\left(\mathcal{M}Q_{n}(i_{n})\right)-Q^{*}_{w^{*}}(i_{n},a_{n})\right].$
Without loss of generality, let $J_{n}(i,a)=w_{n}(r^{j}_{ia}+\gamma\,\mathcal{M}Q_{n}(j))+(1-w_{n})(\mathcal{M}Q_{n}(i))-Q^{*}_{w^{*}}(i,a).$ Also $\beta_{0}$, $\Phi_{0}$ is $\mathcal{G}_{0}$ measurable and $\beta_{n}$, $\Phi_{n}$, $J_{n-1}$ are $\mathcal{G}_{n}$ measurable for $n\geq 1$. Define $J^{Q}_{n}=w^{*}(r^{j}_{ia}+\gamma\mathcal{M}Q_{n}(j))+(1-w^{*})(\mathcal{M}Q_{n}(i))$. We now consider,
$\left|E[J_{n}(i,a)\mid\mathcal{G}_{n}]\right|$
$=\left|E\left[J_{n}(i,a)+J^{Q}_{n}-J^{Q}_{n}+Q^{*}_{w^{*}}(i,a)-Q^{*}_{w^{*}}(i,a)\mid\mathcal{G}_{n}\right]\right|$
$\leq\big{|}\sum_{j=1}^{|S|}p^{j}_{ia}\left(w^{*}\left(r^{j}_{ia}+\gamma\mathcal{M}Q_{n}(j)\right)\right.$
$\quad\left.+(1-w^{*})\mathcal{M}Q_{n}(i)\right)-Q^{*}_{w^{*}}(i,a)\big{|}$
$\quad+\left|w_{n}-w^{*}\right|\left|E\left[r^{j}_{ia}+\gamma\mathcal{M}Q_{n}(j)+\mathcal{M}Q_{n}(i)\mid\mathcal{G}_{n}\right]\right|$
@@
$\quad+\left|w_{n}-w^{*}\right|\left(R_{\max}+2\|Q_{n}\|\right)$
$\leq(1-w^{*}+w^{*}\gamma)\|\Phi_{n}\|+\delta_{n},$
@@
$=E\left[\left(J_{n}(i,a)-E[J_{n}(i,a)|\mathcal{G}_{n}]\right)^{2}|\mathcal{G}_{n}\right]$
$\leq E\left[\left(w_{n}(r^{j}_{ia}+\gamma\mathcal{M}Q_{n}(j))+(1-w_{n})(\mathcal{M}Q_{n}(i))\right)^{2}|\mathcal{G}_{n}\right]$
@@
where $C=\max\{3\kappa R_{\max}^{2}+6(\kappa\gamma^{2}+\kappa_{1})\|Q^{*}_{w^{*}}\|^{2},6(\kappa\gamma^{2}+\kappa_{1})\}$, $w_{n}^{2}<\kappa$ and $(1-w_{n})^{2}<\kappa_{1}$. Therefore, all the conditions of Lemma 1 holds and hence $\Phi_{n}\rightarrow 0$ w.p.1, therefore, $Q_{n}\rightarrow Q^{*}_{w^{*}}$ w.p.1. ∎
@@
Consider the updates $\widetilde{Q}^{A}_{n}$ and $\widetilde{Q}^{B}_{n}$ as in Remark 2. Then $E[V^{BA}_{n}(i,a)|\mathcal{G}_{n}]$ converges to zero w.p.1 as $n\rightarrow\infty$.
@@
$=\widetilde{Q}^{B}_{n}(i,a)-\big{(}(1-\beta_{n})\widetilde{Q}^{A}_{n}(i,a)+\beta_{n}(w(r^{j}_{ia}+\gamma\widetilde{Q}^{B}_{n}(j,b^{*}))$
$\hskip 28.45274pt+(1-w)\widetilde{Q}^{B}_{n}(i,c^{*}))\big{)}$
@@
$V^{BA}_{n+1}(i,a)$ $=(1-\beta_{n})\widetilde{Q}^{B}_{n}(i,a)+\beta_{n}\big{(}w(r^{j}_{ia}+\gamma\widetilde{Q}^{A}_{n}(j,d^{*}))$
$\hskip 28.45274pt+(1-w)\widetilde{Q}^{A}_{n}(i,e^{*})\big{)}-\widetilde{Q}^{A}_{n}(i,a)$
=$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5muii\mkern-1.5mu)}}{{=}}(1-\beta_{n})V^{BA}_{n}(i,a)+\beta_{n}\big{(}w(r^{j}_{ia}+\gamma\widetilde{Q}^{A}_{n}(j,d^{*}))$
$\hskip 28.45274pt+(1-w)\widetilde{Q}^{A}_{n}(i,e^{*}))-\widetilde{Q}^{A}_{n}(i,a)\big{)}.$
@@
$V^{BA}_{n+1}(i,a)$ $=(1-\beta_{n})V^{BA}_{n}(i,a)+\beta_{n}J^{A}_{n}(i,a),$
$V^{BA}_{n+1}(i,a)$ $=(1-\beta_{n})V^{BA}_{n}(i,a)+\beta_{n}J^{B}_{n}(i,a),$
@@
$J^{A}_{n}(i,a)=\widetilde{Q}^{B}_{n}(i,a)-w(r^{j}_{ia}+\gamma\widetilde{Q}^{B}_{n}(j,b^{*}))-(1-w)\widetilde{Q}^{B}_{n}(i,c^{*}),$
$J^{B}_{n}(i,a)=w(r^{j}_{ia}+\gamma\widetilde{Q}^{A}_{n}(j,d^{*}))+(1-w)\widetilde{Q}^{A}_{n}(i,e^{*}))-\widetilde{Q}^{A}_{n}(i,a).$
Therefore,
$E[V^{BA}_{n+1}(i,a)|\mathcal{G}_{n}]$
$=(1-\dfrac{\beta_{n}}{2})E[V^{BA}_{n}(i,a)|\mathcal{G}_{n}]+\dfrac{\beta_{n}}{2}E[w\gamma(\widetilde{Q}^{A}_{n}(j,d^{*})$
$\hskip 28.45274pt-\widetilde{Q}^{B}_{n}(j,b^{*}))+(1-w)(\widetilde{Q}^{A}_{n}(i,e^{*})-\widetilde{Q}^{B}_{n}(i,c^{*}))|\mathcal{G}_{n}].$
@@
$E[V^{BA}_{n+1}(i,a)|\mathcal{G}_{n}]$ $=(1-\zeta_{n})E[V^{BA}_{n}(i,a)|\mathcal{G}_{n}]+\zeta_{n}J^{BA}_{n}(i,a),$ (7)
@@
$J^{BA}_{n}(i,a)$ $=E[w\gamma(\widetilde{Q}^{A}_{n}(j,d^{*})-\widetilde{Q}^{B}_{n}(j,b^{*}))$
$+(1-w)(\widetilde{Q}^{A}_{n}(i,e^{*})-\widetilde{Q}^{B}_{n}(i,c^{*}))|\mathcal{G}_{n}].$
Now for the above recursive Equation (7), we verify the following conditions : (a) $\|E[J^{BA}_{n}|\mathcal{G}_{n}]\|\leq(1-w+w\gamma)\|V^{BA}_{n}\|$ and (b) $Var[J^{BA}_{n}|\mathcal{G}_{n}]\leq C(1+\|V^{BA}_{n}\|)^{2}$. Note that once we verify (a) and (b), then from Lemma 1, we can conclude that the iterative scheme obtained in (7) converges to zero w.p.1. Using the law of total expectation, the condition (b) is trivially satisfied. We now consider,
$|E[J^{BA}_{n}(i,a)|\mathcal{G}_{n}]|$
$=|E[w\gamma(\widetilde{Q}^{A}_{n}(j,d^{*})-\widetilde{Q}^{B}_{n}(j,b^{*}))$
$\hskip 28.45274pt+(1-w)(\widetilde{Q}^{A}_{n}(i,e^{*})-\widetilde{Q}^{B}_{n}(i,c^{*}))|\mathcal{G}_{n}]|$
$=|\sum_{j=1}^{|S|}p^{j}_{ia}w\gamma(\widetilde{Q}^{A}_{n}(j,d^{*})-\widetilde{Q}^{B}_{n}(j,b^{*}))$
$\hskip 28.45274pt+(1-w)(\widetilde{Q}^{A}_{n}(i,e^{*})-\widetilde{Q}^{B}_{n}(i,c^{*}))|$
$=|w\gamma\sum_{j=1,j\neq i}^{|S|}p^{j}_{ia}(\widetilde{Q}^{A}_{n}(j,d^{*})-\widetilde{Q}^{B}_{n}(j,b^{*}))$
$\hskip 28.45274pt+(w\gamma p^{i}_{ia}+(1-w))(\widetilde{Q}^{A}_{n}(i,e^{*})-\widetilde{Q}^{B}_{n}(i,c^{*}))|$
@@
$\hskip 28.45274pt+(w\gamma p^{i}_{ia}+(1-w))|(\widetilde{Q}^{A}_{n}(i,e^{*})-\widetilde{Q}^{B}_{n}(i,c^{*}))|.$
@@
$=\widetilde{Q}^{A}_{n}(j,d^{*})-\widetilde{Q}^{B}_{n}(j,b^{*})\leq\widetilde{Q}^{A}_{n}(j,b^{*})-\widetilde{Q}^{B}_{n}(j,b^{*})\leq\|V^{BA}_{n}\|.$
Now, assume $\widetilde{Q}^{B}_{n}(j,b^{*})>\widetilde{Q}^{A}_{n}(j,d^{*})$. In which case,
$|\widetilde{Q}^{A}_{n}(j,d^{*})-\widetilde{Q}^{B}_{n}(j,b^{*})|$
$=\widetilde{Q}^{B}_{n}(j,b^{*})-\widetilde{Q}^{A}_{n}(j,d^{*})\leq\widetilde{Q}^{B}_{n}(j,d^{*})-\widetilde{Q}^{A}_{n}(j,d^{*})\leq\|V^{BA}_{n}\|.$
@@
$J_{n}(i,a)$ $=w(r^{j}_{ia}+\gamma\widetilde{Q}^{B}_{n}(j,b^{*}))+(1-w)\widetilde{Q}^{B}_{n}(i,c^{*})-Q^{*}_{w}(i,a).$
@@
$\hskip 28.45274pt+(1-w)(\widetilde{Q}^{B}_{n}(i,c^{*})-\widetilde{Q}^{A}_{n}(i,c^{*})).$
Note that $\beta_{0}$, $\Phi_{0}$ is $\mathcal{G}_{0}$ measurable and $\beta_{n}$, $\Phi_{n}$, $J_{n-1}$ are $\mathcal{G}_{n}$ measurable for $n\geq 1$. Now,
$\left|E[J_{n}(i,a)|\mathcal{G}_{n}]\right|$
$=\bigg{|}\sum_{j=1}^{|S|}p^{j}_{ia}\big{(}w(r^{j}_{ia}+\gamma\mathcal{M}\widetilde{Q}^{A}_{n}(j))+(1-w)\mathcal{M}\widetilde{Q}^{A}_{n}(i)\big{)}$
$\hskip 28.45274pt-Q^{*}_{w}(i,a)+E\big{[}w\gamma(\widetilde{Q}^{B}_{n}(j,b^{*})-\widetilde{Q}^{A}_{n}(j,b^{*}))$
$\hskip 39.83368pt+(1-w)(\widetilde{Q}^{B}_{n}(i,c^{*})-\widetilde{Q}^{A}_{n}(i,c^{*}))|\mathcal{G}_{n}\big{]}\bigg{|}$
@@
$\hskip 39.83368pt+w\gamma\left|E\left[(\widetilde{Q}^{B}_{n}(j,b^{*})-\widetilde{Q}^{A}_{n}(j,b^{*}))|\mathcal{G}_{n}\right]\right|$
$\hskip 28.45274pt+\left|(1-w)\right|\left|E[(\widetilde{Q}^{B}_{n}(i,c^{*})-\widetilde{Q}^{A}_{n}(i,c^{*}))|\mathcal{G}_{n}]\right|$
$\leq(1-w+w\gamma)\,\|\Phi_{n}\|+\delta_{n},$
@@
Consider the update rule $Q^{A}_{n}$ and $Q^{B}_{n}$ of model-free double SOR Q-learning as in Algorithm 1. Then $E[W^{BA}_{n}(i,a)|\mathcal{G}_{n}]$ converges to zero w.p.1 as $n\rightarrow\infty$.
@@
Now, we conclude this section with a theoretical justification for the over-estimation bias of the SOR Q-learning algorithm in comparison to the proposed algorithm.
IV-A Over-estimation bias of SOR Q-learning
In Q-learning at every $n^{th}$ iteration the problem is to evaluate $\max_{a}E[Q_{n}(s_{n+1},a)]$, i.e., for set of random variables $\{X_{i}\}^{d}_{i=1}$, we wish to compute $\max_{i}E[X_{i}]$ [8]. Let $\{\mu_{i}\}^{d}_{i=1}$ be the unbiased estimators of $\{X_{i}\}^{d}_{i=1}$. The estimator is unbiased in the sense that $E[X_{i}]=E[\mu_{i}],\forall i$. Therefore, Bias $=E[\max_{i}\mu_{i}]-\max_{i}E[\mu_{i}]$, but from Jensen’s inequality, we have $E[\max_{i}\mu_{i}]\geq\max_{i}E[\mu_{i}]$, which makes Q-learning to over-estimate. Similarly, the estimator of the SORQL algorithm, which is $w\max_{i}\mu_{i}+(1-w)\max_{j}\zeta_{j}$, where $\zeta_{j}$ are unbiased estimators with respect to the set of random variables $\{Y_{j}\}^{d}_{j=1}$ corresponding to the actions in the previous state. Then the bias is given by
$\text{Bias}=E[w\max_{i}\mu_{i}+(1-w)\max_{j}\zeta_{j}]-\max_{i}E[\mu_{i}].$ (8)

### v2

@@ IV Theoretical Analysis
Suppose (A1) holds. Given an MDP defined as in Section II, and $w_{n}$ as in (8). Let every state-action pair be sampled indefinitely. Then, for sample $\{i,a,j,r^{j}_{ia}\}$ the update rule of model-free SOR Q-learning algorithm given by
@@
The correspondence to Lemma 1 follows from associating $X$ with the set of state-action pairs $(i,a)$, $\beta_{n}(x)$ with $\beta_{n}(i,a)$, and $\Phi_{n}(i,a)$ with $Q_{n}(i,a)-Q^{*}_{w^{*}}(i,a)$, where $Q^{*}_{w^{*}}$ is the fixed point of $U_{w^{*}}$. Let the filtration for this process be defined by, $\mathcal{G}_{n}=\{Q_{0},i_{j},a_{j},\beta_{j},w_{j},\forall j\leq n,n\geq 0\}$. Note that $\mathcal{G}_{n}\subseteq\mathcal{G}_{n+1},\forall n\geq 0.$ The iteration of model-free SORQL is rewritten as follows
$\Phi_{n+1}(i_{n},a_{n})$ $=(1-\beta_{n}(i_{n},a_{n}))\Phi_{n}(i_{n},a_{n})$ (15)
@@
$\quad\quad\left.+(1-w_{n})\left(\mathcal{M}Q_{n}(i_{n})\right)-Q^{*}_{w^{*}}(i_{n},a_{n})\right].$
Without loss of generality, let $J_{n}(i,a)=w_{n}(r^{j}_{ia}+\gamma\,\mathcal{M}Q_{n}(j))+(1-w_{n})(\mathcal{M}Q_{n}(i))-Q^{*}_{w^{*}}(i,a).$ Also $\beta_{0}$, $\Phi_{0}$ is $\mathcal{G}_{0}$ measurable and $\beta_{n}$, $\Phi_{n}$, $J_{n-1}$ are $\mathcal{G}_{n}$ measurable for $n\geq 1$. Define $J^{Q}_{n}=w^{*}(r^{j}_{ia}+\gamma\mathcal{M}Q_{n}(j))+(1-w^{*})(\mathcal{M}Q_{n}(i))$. We now consider
$\left|\mathbb{E}[J_{n}(i,a)\mid\mathcal{G}_{n}]\right|$
$=\left|\mathbb{E}\left[J_{n}(i,a)+J^{Q}_{n}-J^{Q}_{n}+Q^{*}_{w^{*}}(i,a)-Q^{*}_{w^{*}}(i,a)\mid\mathcal{G}_{n}\right]\right|$
@@
$\quad\quad\quad\quad+\left|w_{n}-w^{*}\right|\left(R_{\max}+2\|Q_{n}\|\right)$
$\leq(1-w^{*}+w^{*}\gamma)\|\Phi_{n}\|+\delta_{n}$ (16)
@@
$=\mathbb{E}\left[\left(J_{n}(i,a)-\mathbb{E}[J_{n}(i,a)|\mathcal{G}_{n}]\right)^{2}|\mathcal{G}_{n}\right]$
$\leq\mathbb{E}\left[\left(w_{n}(r^{j}_{ia}+\gamma\mathcal{M}Q_{n}(j))+(1-w_{n})(\mathcal{M}Q_{n}(i))\right)^{2}|\mathcal{G}_{n}\right]$
@@
where $C=\max\{3\kappa R_{\max}^{2}+6(\kappa\gamma^{2}+\kappa_{1})\|Q^{*}_{w^{*}}\|^{2},6(\kappa\gamma^{2}+\kappa_{1})\}$, $w_{n}^{2}<\kappa$ and $(1-w_{n})^{2}<\kappa_{1}$.
Consequently, all the conditions of Lemma 1 holds and hence $\Phi_{n}\rightarrow 0$ w.p.1, therefore, $Q_{n}\rightarrow Q^{*}_{w^{*}}$ w.p.1. ∎
@@
Consider the updates $\widetilde{Q}^{A}_{n}$ and $\widetilde{Q}^{B}_{n}$ as in Remark 2. Then $\mathbb{E}[V^{BA}_{n}(i,a)|\mathcal{G}_{n}]$ converges to zero w.p.1 as $n\rightarrow\infty$.
@@
$V^{BA}_{n+1}(i,a)$ =$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5muii\mkern-1.5mu)}}{{=}}(1-\beta_{n})V^{BA}_{n}(i,a)+\beta_{n}\big{(}w(r^{j}_{ia}+\gamma\widetilde{Q}^{A}_{n}(j,d^{*}))$
$\hskip 28.45274pt+(1-w)\widetilde{Q}^{A}_{n}(i,e^{*}))-\widetilde{Q}^{A}_{n}(i,a)\big{)}.$ (18)
@@
$V^{BA}_{n+1}(i,a)$ $=(1-\beta_{n})V^{BA}_{n}(i,a)+\beta_{n}J^{A}_{n}(i,a)$ (19)
$V^{BA}_{n+1}(i,a)$ $=(1-\beta_{n})V^{BA}_{n}(i,a)+\beta_{n}J^{B}_{n}(i,a)$ (20)
@@
$J^{A}_{n}(i,a)=\widetilde{Q}^{B}_{n}(i,a)-w(r^{j}_{ia}+\gamma\widetilde{Q}^{B}_{n}(j,b^{*}))-(1-w)\widetilde{Q}^{B}_{n}(i,c^{*})$ (21)
$J^{B}_{n}(i,a)=w(r^{j}_{ia}+\gamma\widetilde{Q}^{A}_{n}(j,d^{*}))+(1-w)\widetilde{Q}^{A}_{n}(i,e^{*}))-\widetilde{Q}^{A}_{n}(i,a).$ (22)
Therefore
$\mathbb{E}[V^{BA}_{n+1}(i,a)|\mathcal{G}_{n}]$
$=(1-\dfrac{\beta_{n}}{2})\mathbb{E}[V^{BA}_{n}(i,a)|\mathcal{G}_{n}]+\dfrac{\beta_{n}}{2}\mathbb{E}[w\gamma(\widetilde{Q}^{A}_{n}(j,d^{*})$
$\hskip 2.84544pt-\widetilde{Q}^{B}_{n}(j,b^{*}))+(1-w)(\widetilde{Q}^{A}_{n}(i,e^{*})-\widetilde{Q}^{B}_{n}(i,c^{*}))|\mathcal{G}_{n}].$ (23)
@@
$\mathbb{E}[V^{BA}_{n+1}(i,a)|\mathcal{G}_{n}]$ $=(1-\zeta_{n})\mathbb{E}[V^{BA}_{n}(i,a)|\mathcal{G}_{n}]+\zeta_{n}J^{BA}_{n}(i,a)$ (24)
@@
$J^{BA}_{n}(i,a)$ $=\mathbb{E}[w\gamma(\widetilde{Q}^{A}_{n}(j,d^{*})-\widetilde{Q}^{B}_{n}(j,b^{*}))$
$\quad\quad+(1-w)(\widetilde{Q}^{A}_{n}(i,e^{*})-\widetilde{Q}^{B}_{n}(i,c^{*}))|\mathcal{G}_{n}].$ (25)
Now for the above recursive Equation (24), we verify the following conditions : (a) $\|\mathbb{E}[J^{BA}_{n}|\mathcal{G}_{n}]\|\leq(1-w+w\gamma)\|V^{BA}_{n}\|$ and (b) $Var[J^{BA}_{n}|\mathcal{G}_{n}]\leq C(1+\|V^{BA}_{n}\|)^{2}$. Note that once we verify (a) and (b), then from Lemma 1, we can conclude that the iterative scheme obtained in Equation (24) converges to zero w.p.1. Using the law of total expectation, the condition (b) is trivially satisfied. We now consider
$|\mathbb{E}[J^{BA}_{n}(i,a)|\mathcal{G}_{n}]|$
$=|\mathbb{E}[w\gamma(\widetilde{Q}^{A}_{n}(j,d^{*})-\widetilde{Q}^{B}_{n}(j,b^{*}))$
$\hskip 28.45274pt+(1-w)(\widetilde{Q}^{A}_{n}(i,e^{*})-\widetilde{Q}^{B}_{n}(i,c^{*}))|\mathcal{G}_{n}]|$ (26)
@@
$\hskip 28.45274pt+(w\gamma p^{i}_{ia}+(1-w))|(\widetilde{Q}^{A}_{n}(i,e^{*})-\widetilde{Q}^{B}_{n}(i,c^{*}))|.$ (27)
@@
$=\widetilde{Q}^{A}_{n}(j,d^{*})-\widetilde{Q}^{B}_{n}(j,b^{*})\leq\widetilde{Q}^{A}_{n}(j,b^{*})-\widetilde{Q}^{B}_{n}(j,b^{*})\leq\|V^{BA}_{n}\|.$ (28)
Similarly, when $\widetilde{Q}^{B}_{n}(j,b^{*})>\widetilde{Q}^{A}_{n}(j,d^{*})$. We have
$|\widetilde{Q}^{A}_{n}(j,d^{*})-\widetilde{Q}^{B}_{n}(j,b^{*})|$ $\leq\widetilde{Q}^{B}_{n}(j,d^{*})-\widetilde{Q}^{A}_{n}(j,d^{*})\leq\|V^{BA}_{n}\|.$ (29)
@@
$J_{n}(i,a)$
$=w(r^{j}_{ia}+\gamma\widetilde{Q}^{B}_{n}(j,b^{*}))+(1-w)\widetilde{Q}^{B}_{n}(i,c^{*})-Q^{*}_{w}(i,a)$
@@
$\hskip 28.45274pt+(1-w)(\widetilde{Q}^{B}_{n}(i,c^{*})-\widetilde{Q}^{A}_{n}(i,c^{*})).$ (33)
Note that $\beta_{0}$, $\Phi_{0}$ is $\mathcal{G}_{0}$ measurable and $\beta_{n}$, $\Phi_{n}$, $J_{n-1}$ are $\mathcal{G}_{n}$ measurable for $n\geq 1$. Now
$\left|\mathbb{E}[J_{n}(i,a)|\mathcal{G}_{n}]\right|$
@@
$\hskip 28.45274pt+w\gamma\left|\mathbb{E}\left[(\widetilde{Q}^{B}_{n}(j,b^{*})-\widetilde{Q}^{A}_{n}(j,b^{*}))|\mathcal{G}_{n}\right]\right|$
$\hskip 39.83368pt+\left|(1-w)\right|\left|\mathbb{E}[(\widetilde{Q}^{B}_{n}(i,c^{*})-\widetilde{Q}^{A}_{n}(i,c^{*}))|\mathcal{G}_{n}]\right|$
$\leq(1-w+w\gamma)\,\|\Phi_{n}\|+\delta_{n}$ (34)
@@
Consider the update rule $Q^{A}_{n}$ and $Q^{B}_{n}$ of model-free double SOR Q-learning as in Algorithm 1. Then $\mathbb{E}[W^{BA}_{n}(i,a)|\mathcal{G}_{n}]$ converges to zero w.p.1 as $n\rightarrow\infty$.
@@
Now, we conclude this section with a theoretical and experimental analysis of the overestimation bias of the SOR Q-learning algorithm in comparison to the proposed algorithm.
IV-A Theoretical Analysis of Over-estimation Bias in SORQL
In Q-learning at every $n^{th}$ iteration the problem is to evaluate $\max_{a}\mathbb{E}[Q_{n}(s_{n+1},a)]$, i.e., for set of random variables $\{X_{i}\}^{d}_{i=1}$, we wish to compute $\max_{i}\mathbb{E}[X_{i}]$ [11]. Let $\{\mu_{i}\}^{d}_{i=1}$ be the unbiased estimators of $\{X_{i}\}^{d}_{i=1}$. The estimator is unbiased in the sense that $\mathbb{E}[X_{i}]=\mathbb{E}[\mu_{i}],\forall i$. Therefore, Bias $=\mathbb{E}[\max_{i}\mu_{i}]-\max_{i}\mathbb{E}[\mu_{i}]$, but from Jensen’s inequality, we have $\mathbb{E}[\max_{i}\mu_{i}]\geq\max_{i}\mathbb{E}[\mu_{i}]$, which makes Q-learning to over-estimate. Similarly, the estimator of the SORQL algorithm, which is $w\max_{i}\mu_{i}+(1-w)\max_{j}\zeta_{j}$, where $\zeta_{j}$ are unbiased estimators with respect to the set of random variables $\{Y_{j}\}^{d}_{j=1}$ corresponding to the actions in the previous state. Then the bias is given by
$\text{Bias}=\mathbb{E}[w\max_{i}\mu_{i}+(1-w)\max_{j}\zeta_{j}]-\max_{i}\mathbb{E}[\mu_{i}].$ (35)